# Identification and Optimal Nonlinear Control of Turbojet Engine Using Koopman Eigenfunction Model

### v2

@@ 4.4 Autonomous trajectory projection and the complete KEM
Regularization helps excluding non-active modes and preserve only the prominent ones leading to a further dimensionality reduction. The final model is given as
@@
where the RHS can be further approximated using a suitable basis, e.g. polynomials or RBF, or converted to a look-up-table format.
An example of an affine LPV decomposition of the input dynamics $\nabla\mathbf{\Phi}(N)g(N)$ is presented in Appendix A. Finally, the discrete-time matrices can be obtained as $\mathbf{L}=\text{exp}({\mathbf{\Lambda}t})$ and $\mathbf{G_{d}=\Lambda^{-1}(L-I)\nabla\Phi}g$.
@@ 5.1 Proportional-integral controller
[…]
@@ 5.2 LPV-PI controller
The SINDy governing equation was linearized analytically using the following equations:
$a(N)=\frac{\partial f(N)}{\partial N}+\frac{\partial g(N)}{\partial N}W_{f},$ (54)
@@
where $K_{e}=-b(N)/a(N)$ and $T_{e}=-1/a(N)$ are the engine gain and time constant, respectively.
Subsequently, a finite set of sample steady-state points was selected. At each of these points, the system was assumed to be LTI, and the PI controller was tuned to reach optimal performance. There are several methods for LTI PID tuning, e.g. the Skogestad’s internal model method [55] accounting for time delays, or the popular, yet not always sufficient, Ziegler-Nichols methods [56]. In this paper, the MATLAB native command pidtune was utilized. The primary objective was set to reference tracking. Further fine tuning was performed manually via the Live Editor application ”Tune PID Controller” with focus on overshoot reduction. For each LTI subsystem, the trade-off between overshoots and speed was of interest making the process rather time-consuming. Finally, the controller gains were fitted by 5th-order polynomials. It is important to note that these gains are actually the corrected parameters of $K_{p}$ and $K_{i}$. Their relationship is given as
$K_{p}=K_{p,corr}\frac{101325}{p_{1t}}$ (57)
$K_{i}=K_{i,corr}\sqrt{\frac{288}{T_{1t}}}\frac{p_{1t}}{101325}$ (58)
@@ 5.3 SINDy - IMC
Since the dead-time effects are neglected in this study, the standard internal model controller (IMC) approach was considered. Details can be found, e.g., in [57, 58]. The derived SINDy model can be utilized for design of a gain scheduled IMC controller, which is another contribution of this paper. Assuming the fuel system delay as a first-order lag, the second-order internal model in each steady-state point is given as
@@
This expression can be rewritten to an alternative form for $\dot{v}$ control. Multiplying both sides by $K_{e}(\tau_{f}^{2}s^{2}+2\tau_{f}s)E(s)$, performing inverse Laplace transform and integrating both sides yields the following control law:
[…]
where the first term on the right-hand side represents a PID controller with the gains given as $K_{p}=(T_{e}+T_{f})/(K_{e}\tau_{f}^{2})$, $K_{i}=1/(K_{e}\tau_{f}^{2})$ and $K_{d}=T_{e}T_{f}/(K_{e}\tau_{f}^{2})$.
The filter time constant $\tau_{f}$ is the only tuning parameter and can be scheduled as a function of $N$. In this paper, it was assumed constant. The higher the $\tau_{f}$, the lower the sensitivity to measurement noise due to the derivative action as the cut-off frequency decreases, but the slower the response. Thus, the optimal tuning is a trade off between the noise sensitivity and performance. The corrected parameters are assumed as in the previous case. The block diagram of the IMC controller is depicted in Fig. 3.
@@ 5.4 LQI and Kalman observer design
The LQG controller with integral action (servo LQG or LQGI) can be designed in the eigenfunction space to reach a globally optimal control performance even outside the region of validity of the linearized LPV model. The LQGI design concise of two step - the LQI controller design and the Kalman observer design. Both steps can be performed individually according to the separation theorem.
The advantage of designing controllers in the lifted state space, eigenfunction coordinates, is that we can employ powerful tools known from the control theory of linear systems and conveniently analyze the stability and robustness. Opposed to the locally linearized systems, the stability and robustness hold not only in local regions around steady-state operating points, but also in the wider operating region of the engine, supposed that this region is sufficiently covered by training data.
@@ 5.4.1 LQI controller design preliminaries
To design an LQI controller, the eigenfunctions are first augmented with the fuel system dynamics characterized by an FOPDT with time constant $T_{f}$. The first augmented state vector is denoted as $\beta=\begin{bmatrix}\mathbf{\Phi}&W_{f}\end{bmatrix}^{\textbf{T}}$. Since the functions of input dynamics are dependent on the spool speed estimate, the system is evaluated at a finite set of evenly distributed sample points from the $N$ interval. At each point, the system is treated as an LTI system, where the matrices are defined as
@@
A new input variable $v$ is introduced. It represents the fuel flow command given by the controller. Furthermore, $\beta$ is augmented with an integral state $\eta$, dynamics of which are given as
[…]
The final augmented state vector is denoted as $\textbf{z}=\begin{bmatrix}\beta&\eta\end{bmatrix}^{\textbf{T}}$, and the resulting state-space model is
[…]
@@
[…]
This problem is usually solved using quadratic programming. The final K-LQGI control law is defined for corrected quantities. The corrected integral state is given as
[…]
This way, the effects of changing flight conditions are captured to some extent. However, for higher Mach number of flight, the corrections start losing validity due to compressibility effects.
@@ 5.4.2 Optimization algorithm
The metaheuristic approach, e.g. genetic algorithm (GA) or PSO, was utilized to optimize the LQI controller. The MATLAB Riccati equation solver provides a simple way to solve the LQI problem described above. Additionally, alongside with the standard integral cost function (69) minimized by the solver, the GA minimizes a custom cost function based on the integral of absolute error (IAE). Furthermore, since specified stability margins must be respected, the problem is constrained. Thus, GA and PSO are suitable tools [53, 59].
The controller cost function is a sum of a weighted IAE and additional terms accounting for the maximum overshoot and undershoot. These terms also penalize not reaching the maximum and minimum commands fast enough. The weighting by $w_{e}$ is performed so that the error during smaller transients has the same magnitude as during the large transients to even out their importance.
@@
$\min_{\mathbf{Q_{c}}}\sum_{k=1}^{n_{s}-1}w_{e}\left(\frac{|e_{k}|+|e_{k+1}|}{2}\right)\Delta t+\alpha_{os}|\max(N)-1|+\alpha_{us}|\min(N)-0.5|,$ (71)
@@
$e_{k}=N_{d}(t_{k})-N(t_{k}),$ (72)
@@
$\dot{\eta}=0\quad\text{when}\quad v=v_{min}\lor v=v_{max},\quad(\text{Clamping})$ (74)
$W_{f}=\left(1-\Delta t/T_{f}\right)W_{f}+\left(\Delta t/T_{f}\right)v,$ (75)
$\mathbf{\dot{\Phi}=\Lambda\Phi+G}(N)W_{f},$ (76)
$N(t_{k})=\mathbf{C\Phi}(t_{k}),$ (77)
@@
Among the optimized parameters are the weight matrix of the transient spool speed $\mathbf{Q_{N}}$, integral action $\mathbf{Q_{i}}$, delayed fuel flow $\mathbf{Q_{f}}$, and input $\mathbf{R_{c}}$. These are further complemented by a weight perturbation diagonal matrix $\textbf{dQ}_{\varphi}$ targeting individual spectral components. This allows the algorithm to balance a possibly more aggressive integral action. The complete state weighting matrix is given as
$\mathbf{Q_{c}}=\begin{bmatrix}\mathbf{Q_{\varphi}+dQ_{\varphi}}&\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{Q}_{\textbf{f}}&\textbf{0}\\ \textbf{0}&\textbf{0}&\textbf{Q}_{\textbf{i}}\end{bmatrix},$ (81)
where the $\mathbf{Q_{\varphi}}$ matrix is generally obtained as
$\mathbf{x^{T}Q_{x}x}$ $=\mathbf{(C\Phi)^{T}Q_{x}C\Phi}$
$=\mathbf{\Phi^{T}C^{T}Q_{x}C\Phi}$
$=\mathbf{\Phi^{T}Q_{\varphi}\Phi}.$ (82)
Thus, for the GTE, it yields $\mathbf{Q_{\varphi}=C^{T}Q_{N}C}$.
The objective function includes a solution of the continuous algebraic Riccati equation (CARE) for a finite set of sample points $N_{i}$ evenly distributed across the normalization interval $[N_{min},N_{max}]$. At each point, the system is considered LTI with a controllable pair $(\mathbf{A_{aug,i},B_{aug,i}})$, and the following equations are solved:
$\mathbf{K_{LQI,i}=R_{c}^{-1}B_{aug,i}^{T}P_{c}},$ (83)
where $\mathbf{P_{c}}$ is a solution of the CARE given by
$\mathbf{A_{aug,i}^{T}P_{c}+P_{c}A_{aug,i}+Q_{c}-P_{c}B_{aug,i}R_{c}^{-1}B_{aug,i}^{T}P_{c}=0}.$ (84)
The LQI control law obtained from the CARE is given as
$v=-\mathbf{K_{LQI}}(\hat{N})\cdot\begin{bmatrix}\beta&\eta\end{bmatrix}^{\textbf{T}}.$ (85)
Subsequently, the feedback gain matrix entries are interpolated or approximated by a suitable function of $N$, and the numerical integration is utilized to simulate the system’s response to a predefined testing spool-speed command profile depicted in Fig. 4. This profile captures both the small and large transients.
@@ 5.4.3 Kalman observer design
To improve the prediction performance of the baseline KEM, a Kalman observer can be designed. The KEM is considered noisy as follow:
@@
where $\mathbf{w_{p}}$ and $w_{n}$ are the process and measurement noise, respectively, both sampled from zero-mean Normal distribution.
Since both $\mathbf{\Lambda}$ and C matrices are time-invariant, the steady Kalman filter can be obtained as a solution to the estimator CARE. This was discussed in, e.g., [60]. The observer state-space model is shown in Fig. 5.
@@
$\dot{\hat{\beta}}=\textbf{A}_{\textbf{i}}\hat{\beta}+\textbf{B}_{\textbf{i}}v+\textbf{L}_{\beta}(N-\textbf{C}_{\textbf{i}}\hat{\beta})=(\textbf{A}_{\textbf{i}}-\textbf{L}_{\beta}\textbf{C}_{\textbf{i}})\hat{\beta}+\textbf{B}_{\textbf{i}}v+\textbf{L}_{\beta}N,$ (87)
where the Kalman gain matrix $\textbf{L}_{\beta}$ is considered constant.
As can be seen, the fuel flow state $W_{f}$ is also included, since the equation is formulated for $\beta$. The gain matrix can be further written as $\textbf{L}_{\beta}=\begin{bmatrix}\textbf{L}&0\end{bmatrix}^{\textbf{T}}$, where, during the design phase, the matrix L is optimized only for $\mathbf{\Phi}$, yielding a static Kalman observer, and the fuel flow estimation depends only on the first-order model. Thus, $\hat{W}_{f}=W_{f}$ as demonstrated in Eq. (88).
$\hat{W}_{f}=W_{f}=\left(1-\frac{\Delta t}{T_{f}}\right)\hat{W}_{f}+\frac{\Delta t}{T_{f}}v.$ (88)
Thus, the $W_{f}$ dynamics are accounted for only during the LQI design. This assumption was deemed valid since the first-order model of the fuel system was considered accurate enough to give reliable information about the fuel flow with no need for correction step.
If the estimation of $W_{f}$ would still be of interest, it would be more advantageous to switch to the discrete-time formulation of the KEM and utilize the time-varying Kalman filter, as the $\mathbf{A_{i}}$ matrix would become time-varying.
Several approaches can be used for the observer design, see, e.g., [60, 61, 62]. In this paper, similarly to the LQI design step, assuming that the pair $(\mathbf{\Lambda,C})$ is observable, the CARE approach is considered, and the Kalman gain is given by the following equations:s
$\mathbf{L=P_{o}C^{T}R_{o}^{-1}},$ (89)
@@
$\mathbf{\Lambda P_{o}+P_{o}\Lambda^{T}+G_{o}Q_{o}G_{o}^{T}-P_{o}C^{T}R_{o}^{-1}CP_{o}=0},$ (90)
where $\mathbf{G_{o}}$ is the process noise gain matrix usually set to I, $\mathbf{R_{o}}$ is the measurement noise covariance matrix estimated from data, and $\mathbf{Q_{o}}$ is the process noise covariance matrix. The latter can be estimated using the test dataset eigenfunction error covariance $\text{Cov}(\mathbf{\Phi}(t)-\mathbf{\Phi}(N))$, where $\mathbf{\Phi}(N)$ are eigenfunctions evaluated using the ground-truth data. Alternatively, $\mathbf{Q_{o}}$ can be estimated also from the eigenfunction time derivative error covariance since the process noise $\mathbf{w_{p}}$ is acting on the dynamics in Eq. (86). Thus, the process noise covariance estimate is given as
$\mathbf{\Sigma_{p}}\approx\text{Cov}\left(\frac{\partial\mathbf{\Phi}(N)}{\partial N}\dot{N}-\mathbf{\Lambda\Phi}(t)-\nabla\mathbf{\Phi}\big{(}\hat{N}(t)\big{)}g\big{(}\hat{N}(t)\big{)}u(t)\right).$ (91)
@@
$v=-\mathbf{K_{\beta}}(\hat{N})\hat{\beta}-K_{i}(\hat{N})\eta,$ (92)
and $\mathbf{K_{LQI}}=\begin{bmatrix}\textbf{K}_{\beta}&K_{i}\end{bmatrix}$.
Inserting the control law in Eq. (92) to Eq. (67) and augmenting with the Kalman filter, Eq. (87), according to the separation theorem [62], we obtain the following closed-loop state-space model:
$\begin{bmatrix}\dot{\beta}\\ \dot{\varepsilon}\\ \dot{\eta}\end{bmatrix}=\begin{bmatrix}\mathbf{A_{i}-B_{i}K_{\beta}}&\mathbf{B_{i}K_{\beta}}&-\textbf{B}_{\textbf{i}}K_{i}\\ \textbf{0}&\mathbf{A_{i}-L_{\beta}C_{i}}&\textbf{0}\\ -\textbf{C}_{\textbf{i}}&\textbf{0}&0\end{bmatrix}\begin{bmatrix}\beta\\ \varepsilon\\ \eta\end{bmatrix}+\begin{bmatrix}\textbf{0}\\ \textbf{0}\\ 1\end{bmatrix}N_{d},$ (93)
@@
where $\hat{\beta}=\begin{bmatrix}\hat{\mathbf{\Phi}}&\hat{W}_{f}\end{bmatrix}^{\textbf{T}}$ and $\varepsilon=\beta-\hat{\beta}=\begin{bmatrix}\mathbf{\Phi-\hat{\Phi}}&0\end{bmatrix}^{\textbf{T}}$ are the augmented state and estimation error vectors, respectively.
It can be shown that by inserting $\varepsilon$ to Eq. (93), the FOPDT fuel flow model, Eq. (88), will be revealed in the closed-loop structure. In the open-loop model, the $-\textbf{C}_{\textbf{i}}$ matrix in the second row is replaced by 0, as the integral of the tracking error changes to the integral of the demanded output only. Note that $\eta$ is computed using the real measured $N$, not $\hat{N}$, but the latter can be used, e.g., for diagnostics.

### v6

@@ 4.4 Autonomous trajectory projection and the complete KEM
Regularization helps in exclusion of the inactive modes and preserves only the prominent ones, leading to further dimensionality reduction. The final model is given as
@@
where the RHS can be further approximated using a suitable basis, e.g., polynomials or RBF, or converted to a look-up-table format.
Unlike the locally linearized interpolated LPVs, this system is valid across the entire operating region of the GTE. Furthermore, the only nonlinear component is the input dynamics, which can be sampled by evaluation at arbitrary points in the state space. The sampled subsystems are LTI and coherent to the underlying dynamics. This will be leveraged during the control design.
An example of an affine LPV decomposition of the input dynamics $\nabla\mathbf{\Phi}(N)g(N)$ is presented in Appendix A. The discrete-time matrices can be obtained as $\mathbf{L}=\text{exp}({\mathbf{\Lambda}t})$ and $\mathbf{G_{d}=\Lambda^{-1}(L-I)\nabla\Phi}g$.
@@ 5.1 Proportional-integral controller
A control system with a PI controller and fuel-flow limiters was designed using the nonlinear CLM. According to the PART 31 regulation, the control system should be able to safely increase the thrust from 15 $\%$ to 95 $\%$ of the nominal value in under 5 seconds [bib54]. For this purpose, the PI controller was manually tuned to achieve the desired performance.
@@ 5.2 LPV-PI controller
To simplify the control design in the frequency domain, the SINDy governing equation was linearized analytically using the following equations:
$a(N,W_{f})=\frac{\partial f(N)}{\partial N}+\frac{\partial g(N)}{\partial N}W_{f},$ (57)
@@
where $K_{e}=-b/a$ and $T_{e}=-1/a$ are the engine gain and time constant, respectively.
Subsequently, a finite set of sample steady-state points, $N_{i}$ and $W_{f}(N_{i})$, was selected. The steady-state fuel flow was obtained as
$W_{f}(N_{i})=-f(N_{i})/g(N_{i}).$ (60)
At each of these points, the system was assumed to be LTI, and the PI controller was tuned to reach the optimal performance. There are several methods for LTI PID tuning, such as Skogestad’s internal model method [bib55], which accounts for time delays, or the popular yet not always sufficient Ziegler-Nichols methods [bib56]. In this paper, the PI controller was tuned to reach a fast step response with no overshoot, minimizing the integral of quadratic tracking error. To avoid unreasonable gain values, additional regularization of the gains was included. Finally, the controller gains were fitted by 5th-order polynomials.
To handle varying flight conditions, the gains obtained from the design were treated as corrected gains, $K_{p,corr}$ and $K_{i,corr}$. The true gains, $K_{p}$ and $K_{i}$, were computed online as
$K_{p}$ $=K_{p,corr}\bigg(\frac{101325}{p_{1t}}\bigg),$ (61)
$K_{i}$ $=K_{i,corr}\sqrt{\frac{288}{T_{1t}}}\bigg(\frac{p_{1t}}{101325}\bigg),$ (62)
where $p_{1t}$ and $T_{1t}$ were measured in real time.
@@ 5.3 SINDy - IMC
Since the dead-time effects are neglected in this study, the standard internal model controller (IMC) approach was considered. Details can be found, e.g., in [bib57, bib58]. The derived SINDy model can be utilized for the design of a gain-scheduled IMC controller, which is another proposition of this paper. Assuming the fuel system delay as a first-order lag, the second-order internal model in each steady-state point is given as
@@
This expression can be rewritten to an alternative form for $\dot{v}$ control. Multiplying both sides by $K_{e}(\tau_{f}^{2}s^{2}+2\tau_{f}s)E(s)$, performing an inverse Laplace transform, and integrating both sides yields the following control law:
$\dot{v}$ $=\frac{1}{K_{e}\tau_{f}^{2}}\left(T_{e}T_{f}\dot{e}+(T_{e}+T_{f})e+\int_{0}^{t}e(\tau)d\tau\right)$
$-\frac{2}{\tau_{f}}v,$ (67)
$v$ $=\int\dot{v}d\tau,$ (68)
where the first term on the right-hand side represents a PID controller with the corrected gains given as $K_{p,corr}=(T_{e}+T_{f})/(K_{e}\tau_{f}^{2})$, $K_{i,corr}=1/(K_{e}\tau_{f}^{2})$, and $K_{d,corr}=T_{e}T_{f}/(K_{e}\tau_{f}^{2})$.
The filter time constant $\tau_{f}$ is the only tuning parameter and can be scheduled as a function of $N$. In this paper, it was assumed to be constant. The higher the $\tau_{f}$, the lower the sensitivity to measurement noise due to the derivative action as the cut-off frequency decreases, but the slower the response. Thus, the optimal tuning is a trade-off between the noise sensitivity and performance. The corrected parameters are assumed to be as in the previous case. The block diagram of the IMC controller is depicted in Fig. 4.
@@ 5.4 LQI and Kalman observer design
The LQG controller with integral action (servo LQG or LQGI) can be designed in the eigenfunction space to reach a globally optimal control performance even outside the region of validity of the linearized LPV model. The LQGI design consists of two steps: the LQI controller design and the Kalman observer design. Both steps can be performed individually according to the separation theorem.
The advantage of designing controllers in the eigenfunction coordinates is that we can employ powerful tools known from the control theory of linear systems and conveniently analyze the stability and robustness. Opposed to the locally linearized systems, the stability and robustness hold not only in local regions around steady-state operating points, but also in the wider operating region of the engine, assuming that this region is sufficiently covered by training data.
@@ 5.4.1 LQI controller design preliminaries
To design an LQI controller, the eigenfunctions are first augmented with the fuel system dynamics characterized by an FOPDT with time constant $T_{f}$. The first augmented state vector is denoted as $\boldsymbol{\zeta}=\begin{bmatrix}\mathbf{\Phi}&W_{f}\end{bmatrix}^{\intercal}$. Since the functions of input dynamics in the system (4.4) are dependent on the spool speed estimate, the system is evaluated at a finite set of evenly distributed sample points, $N_{i}$, from the $N$ interval. At each point, the system is treated as an LTI system. The fuel system dynamics (10) in state-space form are augmented with the system, and the resulting matrices are defined as
@@
A new input variable $v$ is introduced. It represents the fuel flow command given by the controller. Furthermore, $\boldsymbol{\zeta}$ is augmented with an integral state $\eta$, dynamics of which is given as
$\eta=\int_{t_{0}}^{t_{f}}(N_{d}-N)d\tau,\hskip 14.22636pt\dot{\eta}=N_{d}-\textbf{C}\mathbf{\Phi}.$ (70)
The final augmented state vector is denoted as $\textbf{z}=\begin{bmatrix}\boldsymbol{\zeta}&\eta\end{bmatrix}^{\intercal}$, and the resulting state-space model is
$\dot{\textbf{z}}=\mathbf{A_{aug,i}\hskip 1.42262ptz}+\mathbf{B_{aug,i}}\hskip 1.42262ptv+\begin{bmatrix}\textbf{0}\\ 1\end{bmatrix}N_{d},$ (71)
@@
$J_{LQI}=\int_{t_{0}}^{t_{f}}\left(\textbf{z}^{\intercal}\mathbf{Q_{z}}\textbf{z}+Rv^{2}\right)d\tau,$ (73)
where $\mathbf{Q_{z}}$ and $R$ are the augmented state and input weighting matrices, respectively.
The final K-LQGI control law is defined for normalized corrected quantities. Thus, the integral state in (70) must also be computed by first correcting (and normalizing), and then integrating the tracking error.
This way, the effects of changing flight conditions are captured to some extent. However, for higher Mach numbers of flight, the corrections start losing validity due to compressibility effects.
@@ 5.4.2 Optimization algorithm
MATLAB provides a simple way to solve the LQI problem (73) using the Riccati equation. Another task is to design the weighting matrices such that the overall performance is optimal. For this purpose, the metaheuristic approach, e.g., genetic algorithm (GA) or PSO, can be utilized. The algorithm minimizes a custom cost function based on the integral of absolute error (IAE). Furthermore, since the specified stability margins must be respected, the problem is constrained. Thus, GA and PSO are suitable tools [bib53, bib59].
The global controller cost function is a sum of a weighted IAE and additional terms accounting for the maximum overshoot and undershoot of the maximum value, $\bar{N}_{max}=1$, and the minimum value, $\bar{N}_{min}=0.5$, respectively, where $\bar{N}$ denotes the physical spool speed relativized using the nominal value of 14000 RPM ($\bar{N}=N_{phys}/14000$). These terms also penalize not reaching the maximum and minimum commands fast enough. The weighting by $w_{e}$ is performed so that the error during smaller transients has the same magnitude as during the large transients to even out their importance.
@@
$\min_{\mathbf{Q_{z}},R}$ $\sum_{k=1}^{n_{s}-1}w_{e}\left(\frac{|e_{k}|+|e_{k+1}|}{2}\right)\Delta t$
$+\alpha_{os}|\max(\bar{N})-1|$
$+\alpha_{us}|\min(\bar{N})-0.5|,$ (74)
@@
$e_{k}=\bar{N}_{d}(t_{k})-\bar{N}(t_{k}),$ (75)
$\bar{N}(t_{k})=(10000N(t_{k})+5000)/14000,$ (76)
$N(t_{k})=\mathbf{C\Phi}(t_{k}),$ (77)
$\mathbf{\dot{\Phi}=\Lambda\Phi+G}(N)W_{f},$ (78)
$W_{f,k+1}=\left(1-\Delta t/T_{f}\right)W_{f,k}+\left(\Delta t/T_{f}\right)v_{k},$ (79)
@@
$\dot{\eta}=0\quad\text{when}\quad v=v_{min}\lor v=v_{max},$ (81)
@@
Numerical integration is utilized to simulate the system’s response to a predefined testing spool-speed command profile depicted in Fig. 5. This profile captures both the small and large transients. The response is used for cost calculation according to (5.4.2).
@@
Among the optimized parameters are the weight matrix of the transient spool speed $\mathbf{Q_{N}}$, integral action $\mathbf{Q_{i}}$, delayed fuel flow $\mathbf{Q_{f}}$, and input $\mathbf{R_{c}}=R$. These are further complemented by a weight perturbation diagonal matrix $\mathbf{dQ_{\Phi}}$ targeting individual spectral components. This allows the algorithm to balance a possibly more aggressive integral action. The complete state weighting matrix is given as
$\mathbf{Q_{z}}=\begin{bmatrix}\mathbf{Q_{\Phi}+dQ_{\Phi}}&\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{Q}_{\textbf{f}}&\textbf{0}\\ \textbf{0}&\textbf{0}&\textbf{Q}_{\textbf{i}}\end{bmatrix}$, (85)
where the $\mathbf{Q_{\Phi}}$ matrix is generally obtained as
$\mathbf{x^{\intercal}Q_{x}x}$ $=\mathbf{(C\Phi)^{\intercal}Q_{x}C\Phi}$
$=\mathbf{\Phi^{\intercal}C^{\intercal}Q_{x}C\Phi}$
$=\mathbf{\Phi^{\intercal}Q_{\Phi}\Phi}.$ (86)
Thus, for the GTE, it yields $\mathbf{Q_{\Phi}=C^{\intercal}Q_{N}C}$.
The input $v$ follows a closed-loop LQI control law, which will be presented in the following paragraphs.
As before, the control law is designed for a finite set of sample points $N_{i}$ evenly distributed across the normalization interval $N_{min},N_{max}]$. At each point, the system is considered LTI with a controllable pair $(\mathbf{A_{aug,i},B_{aug,i}})$, and the following equations are solved:
$\mathbf{K_{LQI,i}=R_{c}^{-1}B_{aug,i}^{\intercal}P_{c,i}},$ (87)
where $\mathbf{P_{c,i}}$ is a corresponding solution of the CARE
$\mathbf{A_{aug,i}^{\intercal}P_{c,i}}$ $\mathbf{+P_{c,i}A_{aug,i}+Q_{z}}$
$\mathbf{-P_{c,i}B_{aug,i}R_{c}^{-1}B_{aug,i}^{\intercal}P_{c,i}=0},$ (88)
where $\mathbf{R_{c}}$ is the controller input weighting matrix.
Subsequently, the feedback gain matrix entries are interpolated or approximated by a suitable function of $N$, and the resulting gain-scheduled LQI control law is given as
$v=-\mathbf{K_{LQI}}(\hat{N})\hskip 1.42262pt\begin{bmatrix}\boldsymbol{\zeta}&\eta\end{bmatrix}^{\intercal}$. (89)
@@ 5.4.3 Kalman observer design
To improve the prediction performance of the baseline KEM, a Kalman observer can be designed. Introducing noise, the KEM equations are given as
@@
where $\mathbf{w_{p}}$ and $w_{n}$ are the process and measurement noise, respectively, both assumed to be Gaussian white noise with covariances $\mathbf{\Sigma_{p}}$ and $\sigma_{n}^{2}$, respectively.
Since both $\mathbf{\Lambda}$ and C matrices are time-invariant, the steady Kalman filter can be obtained as a solution to the estimator CARE. This was discussed in, e.g., [bib60]. The observer state-space model is shown in Fig. 6.
@@
[…]
$=(\mathbf{A_{i}}-\textbf{L}_{\boldsymbol{\zeta}}\mathbf{C_{i}})\hat{\boldsymbol{\zeta}}+\mathbf{B_{i}}v+\textbf{L}_{\boldsymbol{\zeta}}N,$ (91)
where the Kalman gain matrix $\textbf{L}_{\boldsymbol{\zeta}}$ is constant.
As can be seen, the fuel flow state $W_{f}$ is also included, since the equation is formulated for $\boldsymbol{\zeta}$. The gain matrix can be further written as $\textbf{L}_{\boldsymbol{\zeta}}=\begin{bmatrix}\textbf{L}&0\end{bmatrix}^{\intercal}$, where, during the design phase, the matrix L is optimized only for $\mathbf{\Phi}$, yielding a static Kalman observer, and the fuel flow estimation depends only on the first-order model. Thus, $\hat{W}_{f}=W_{f}$ as demonstrated in the following equation:
$\hat{W}_{f,k+1}=W_{f,k+1}=\left(1-\frac{\Delta t}{T_{f}}\right)\hat{W}_{f,k}+\frac{\Delta t}{T_{f}}v_{k}.$ (92)
Thus, the $W_{f}$ dynamics are accounted for only during the LQI design. This assumption was deemed valid since the first-order model of the fuel system was considered accurate enough to provide reliable information about the fuel flow with no need for Kalman correction.
If the estimation of $W_{f}$ is still of interest, it would be more advantageous to switch to the discrete-time formulation of the KEM and utilize the time-varying Kalman filter, as the system matrix would become time-varying.
Several approaches can be used for the observer design, see, e.g., [bib60, bib61, bib62]. In this paper, similarly to the LQI design step, assuming that the pair $(\mathbf{\Lambda,C})$ is observable, the CARE approach is considered, and the Kalman gain is given by the following equations:s
$\mathbf{L=P_{o}C^{\intercal}R_{o}^{-1}},$ (93)
@@
$\mathbf{\Lambda P_{o}+P_{o}\Lambda^{\intercal}+G_{o}Q_{o}G_{o}^{\intercal}-P_{o}C^{\intercal}R_{o}^{-1}CP_{o}=0},$ (94)
where $\mathbf{G_{o}}$ is the process noise gain matrix usually set to I, $\mathbf{R_{o}}$ is the measurement noise covariance matrix estimated from data, and $\mathbf{Q_{o}}$ is the process noise covariance matrix. The latter can be estimated using the test dataset eigenfunction dynamics error covariance $\text{Cov}(\Delta\mathbf{\Phi}(t)-\Delta\mathbf{\Phi}(N(t)))/\Delta t$, where $\mathbf{\Phi}(N(t))$ are eigenfunctions evaluated using the ground-truth data. Alternatively, $\mathbf{Q_{o}}$ can also be estimated from the eigenfunction time derivative error covariance since the process noise $\mathbf{w_{p}}$ is acting on the dynamics in (90). Thus, the process noise covariance estimate is given as
$\mathbf{\Sigma_{p}}\approx\text{Cov}\bigg(\frac{\partial\mathbf{\Phi}(N(t))}{\partial N}\dot{N}$ $-\mathbf{\Lambda\Phi}(t)$
$-\nabla\mathbf{\Phi}\big(\hat{N}(t)\big)g\big(\hat{N}(t)\big)u(t)\bigg).$ (95)
@@
$v=-\mathbf{K_{\boldsymbol{\zeta}}}(\hat{N})\hat{\boldsymbol{\zeta}}-K_{i}(\hat{N})\eta,$ (96)
and $\mathbf{K_{LQI}}=\begin{bmatrix}\textbf{K}_{\boldsymbol{\zeta}}&K_{i}\end{bmatrix}$.
Inserting the control law in (96) to (71) and augmenting with the Kalman filter, (5.4.3), according to the separation theorem [bib62], we obtain the following closed-loop state-space model:
$\begin{bmatrix}\dot{\boldsymbol{\zeta}}\\ \dot{\boldsymbol{\varepsilon}}\\ \dot{\eta}\end{bmatrix}$ $=\begin{bmatrix}\mathbf{A_{i}-B_{i}K_{\boldsymbol{\zeta}}}&\mathbf{B_{i}K_{\boldsymbol{\zeta}}}&-\mathbf{B_{i}}K_{i}\\ \textbf{0}&\mathbf{A_{i}-L_{\boldsymbol{\zeta}}C_{i}}&\textbf{0}\\ -\mathbf{C_{i}}&\textbf{0}&0\end{bmatrix}\begin{bmatrix}\boldsymbol{\zeta}\\ \boldsymbol{\varepsilon}\\ \eta\end{bmatrix}$
$+\begin{bmatrix}\textbf{0}\\ \textbf{0}\\ 1\end{bmatrix}N_{d},$ (97)
@@
where $\hat{\boldsymbol{\zeta}}=\begin{bmatrix}\hat{\mathbf{\Phi}}&\hat{W}_{f}\end{bmatrix}^{\intercal}$ and $\boldsymbol{\varepsilon=\zeta-\hat{\zeta}}=\begin{bmatrix}\mathbf{\Phi-\hat{\Phi}}&0\end{bmatrix}^{\intercal}$ are the augmented state and estimation error vectors, respectively.
In the open-loop model, the $\mathbf{C_{i}}$ matrix in the second row is replaced by 0, as the integral of the tracking error changes to the integral of the demanded output only. Note that $\eta$ is computed using the real measured $N$, not $\hat{N}$. Given the noisy spool-speed measurement, the controller gains and KEM input dynamics are evaluated using the Kalman estimate, $\hat{N}$.